# Federated Learning for Smart Grid: A Survey on Applications and Potential Vulnerabilities

### v2

@@ Abstract.
The Smart Grid (SG) is a critical energy infrastructure that collects real-time electricity usage data to forecast future energy demands using information and communication technologies (ICT). Due to growing concerns about data security and privacy in SGs, federated learning (FL) has emerged as a promising training framework. FL offers a balance between privacy, efficiency, and accuracy in SGs by enabling collaborative model training without sharing private data from IoT devices. In this survey, we thoroughly review recent advancements in designing FL-based SG systems across three stages: generation, transmission and distribution, and consumption. Additionally, we explore potential vulnerabilities that may arise when implementing FL in these stages. Finally, we discuss the gap between state-of-the-art FL research and its practical applications in SGs and propose future research directions. These focus on potential attack and defense strategies for FL-based SG systems and the need to build a robust FL-based SG infrastructure. Unlike traditional surveys that address security issues in centralized machine learning methods for SG systems, this survey specifically examines the applications and security concerns in FL-based SG systems for the first time. Our aim is to inspire further research into applications and improvements in the robustness of FL-based SG systems.
@@
copyright: acmlicensedisbn: ccs: Security and privacy Distributed systems security
@@ 1. Introduction
The Smart Grid (SG), which integrates the physical infrastructure of the electrical grid with modern information and communication technologies (ICTs) (vasisht2022smart), is a prime example of a cyber-physical system (CPS). Unlike conventional power grids, SGs can detect, respond to, and proactively manage changes in electricity demand and supply, while also addressing various operational challenges (obeagu2022overview).
The SG architecture incorporates several key components that work together to enhance the overall functionality of the electrical system (mittal2024smart), including advanced metering infrastructure (AMI) (molokomme2019survey), communication networks (hu2023multi), and sophisticated information technology solutions (reza2014overview), among others. By leveraging real-time data and advanced analytics, SGs aim to improve the efficiency of power generation, transmission and distribution, and consumption (chopda2021overview). Although SGs enable a more flexible and resilient power system that can better accommodate the growing demand for electricity, the increasing reliance on ICTs has made SG systems more vulnerable to security and privacy attacks. Federated Learning (FL) has emerged as a promising approach to addressing privacy and security concerns in SG systems. This innovative distributed machine learning paradigm enables collaborative model training across multiple participants without the need to share raw data, making it particularly suitable for SG applications, where generators are distributed and data privacy for each generator is crucial. We categorize the application of FL in SGs into three aspects: generation, transmission and distribution, and consumption. From the perspective of generation (tuballa2016review) in SGs, FL is primarily being used to enhance renewable energy integration and forecasting (camacho2011control). FL enables distributed SG generators to collaboratively train models for predicting resources such as solar and wind power output without sharing sensitive operational data. In the transmission (bayoumi2015power) and distribution aspects (hamidi2010smart) of SGs, FL is proving valuable for line fault detection and classification, optimal power flow (usman2022novel), voltage control (xie2021review), and more. By utilizing a FL-based approach, grid operators can quickly identify and locate faults in transmission lines, while collaboratively training models for power distribution without compromising the privacy of individual generators. In the consumption process of SGs (kaur2020energy), FL is extensively used for load forecasting (husnoo2023secure) and energy theft detection (li2024fuse). Smart meters can collaboratively train models to predict household energy consumption patterns without sharing individual usage data.
@@
Ref Year Stage Applications DL-based Privacy Security FL-based
Grataloup et al. (grataloup2024review) 2024 $\surd$ $\surd$ $\surd$ $\surd$
Zheng et al. (zheng2024advancing) 2024 $\surd$ $\surd$ $\surd$ $\surd$
Zhang et al. (zhang2024vulnerability) 2024 $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
Pandya et al. (pandya2023federated) 2023 $\surd$ $\surd$ $\surd$ $\surd$
Cheng et al. (cheng2022review) 2022 $\surd$ $\surd$ $\surd$ $\surd$
Mirzaee et al. (mirzaee2022smart) 2022 $\surd$ $\surd$ $\surd$
Ding et al. (ding2022cyber) 2022 $\surd$ $\surd$ $\surd$ $\surd$
Cui et al. (cui2020detecting) 2020 $\surd$ $\surd$ $\surd$ $\surd$
@@ 1.1. Related Surveys
We summarize the state-of-the-art (SOTA) survey papers focus on SGs, with particular attention to seven aspects: whether the paper discusses the stages of SGs, compares SG applications, includes Deep Learning (DL)-based methods, discusses privacy and security issues, and whether it focuses on FL-based SG systems (see Table 1).
Grataloup et al. (grataloup2024review) focus on the integration of FL to address challenges in renewable energy applications, such as energy management, system reliability, energy consumption prediction, and load forecasting. However, this survey only highlights FL use cases in SGs as privacy-preserving collaborative learning approaches and overlooks the potential vulnerabilities of SGs when using deep learning and federated learning technologies. Zheng et al. (zheng2024advancing) provide a detailed discussion of components in FL systems, including model aggregation, data partitioning, and local model training. However, their survey paper lacks a discussion on model-related attacks within FL-based SG systems. Zhang et al. (zhang2024vulnerability) offer a comprehensive survey of applications for each stage of SG systems and discuss potential privacy and security violations. However, their focus is primarily on centralized machine learning algorithms. Our work bridges the gaps between SG, FL, and privacy and security vulnerabilities. Pandya et al. (pandya2023federated) demonstrate a broader scope, covering smart cities with a FL framework to preserve privacy. This approach enables effective utilization of large-scale data in areas such as electric supply chain monitoring, air quality forecasting, water consumption prediction, traffic management, and pandemic diagnosis. However, this work lacks details on SG applications and security issues within FL frameworks. Cheng et al. (cheng2022review) conduct a survey of the FL paradigm, comparing the types, network structures, and global model aggregation methods in SG. However, the brief survey lacks a discussion on robustness. Mirzaee et al. (mirzaee2022smart) and Ding et al. (ding2022cyber) discuss conventional security threats to SGs as well as emerging issues related to the application of machine learning techniques. They cover various types of cyberattacks, privacy concerns related to smart meter data, and potential countermeasures to address these challenges. However, they only cover a limited number of applications in SGs and primarily focus on centralized cases. Cui et al. (cui2020detecting) mainly focus on false data attacks and privacy issues in SGs. Compared to previous literature, our paper provides a comprehensive discussion on FL-based SGs and their potential vulnerabilities. We hope this can complement existing research and inspire further studies in this direction.
@@ 1.2. Key Contributions
    Our survey extensively discusses SOTA attack and defense methods in FL and identifies a significant gap between FL research and its applications in FL-based SG systems. We further suggest directions for future exploration in this domain.
@@ 1.3. Structure of the Paper
The structure of our paper is illustrated in Figure 1. In the following sections, we begin by exploring the foundational terminologies of SGs and the FL framework in Section 2. Next, Section 3 provides a comprehensive review of FL-based SG applications, detailing their implementation across various stages. In Section 4, we thoroughly examine the vulnerabilities in SG applications, covering both centralized machine learning and FL-based SG applications. Section 5 presents an overview of commonly employed attack and defense methods within general FL frameworks, identifies existing gaps, and proposes future research directions in the field. Finally, Section 6 concludes the paper.
@@ 2.1. Smart Grid
The electric power grid has evolved by leaps and bounds since its inception (masera2018smart). A major development in this domain was the integration of ICT. The integration of ICT with the traditional power grid facilitated distributed power generation via better resource management and automated control (simoes2012comparison). ICT integration also marked the evolution of the conventional grid into the SG and solved several limitations including unidirectional power flow and centralized bulk generation of electricity. Additionally, it reduced the power and energy domain’s dependence on fossil fuels like coal and natural gas, leading the way for sustainability and decarbonization via the development of microgrids (jahromi2023cyber; yadav2020microgrid). The SG power flow framework consists of the following stages:
@@
    Generation: This stage refers to the production of AC/DC power. In the conventional power grid, power generation was centralized in nature. However, with ICT integration for real-time monitoring, power generation can also be achieved in a regulated manner at residential buildings (for example using solar photovoltaic panels) and used for local consumption/fed back to the grid (kahrobaee2012multiagent).
@@
    Transmission/Distribution: This stage refers to the transfer of the power produced at the generation stage to the appropriate end nodes for utilization. This involves the use of transformers, transmission lines, substation equipment and other line devices like fault detector relays, breakers to ensure a safe and reliable transfer at appropriate bus voltage levels. The fault detector relays and breakers are often digital in nature and controlled by the central monitoring station via ICT devices (gaspar2023smart).
@@
    Consumption: This is final stage of the SG power flow process and involves the utilization of the power as received from the transmission and distribution networks for powering electrical devices. The consumption level can involve entities like government agencies, private industries, private individuals, etc., each of which can have a different consumption footprint. In the modern SG, integration of ICT devices means that some consumer-level devices like electric vehicles can also be used to transfer surplus/unused power back to the grid (coordinated via V2G communications (madawala2011bidirectional)).
@@
While ICT integration in the power grid enabled easier and more robust resource management from a single centralized Supervisory Control and Data Acquisition (SCADA)-based control system (gaushell1993scada), the increasing number of IoT devices and local generation facilities like solar panels means that it can be extremely computationally resource intensive to monitor them. Inadequate monitoring facilities can mean that intrusive events like cyberattacks and line faults may not be detected and mitigated on time.
@@
Slowed mitigation of cyberattacks, faults and other control emergencies can have a devastating impact on the stability of the SG (jahromi2023cyber). Moreover, the SG being a cyber-physical system implies that cyber layer vulnerabilities can be exploited to cause severe physical ramifications. For example, cyberattacks on the SG can lead to uncontrolled voltage increase at one or more buses leading to equipment damage and consumer-level device burnout (rath2022behind). To alleviate these issues, several researchers have proposed, implemented and evaluated the use of Artificial Intelligence (AI)-based tools and techniques for applications like real-time control (bose2017artificial), cybersecurity (ruan2023super), load forecasting (zhou2017holographic), and fault detection (sun2023high). These AI tools are typically deployed at the centralized controller level for monitoring, removing the need for human system operators.
@@
As the AI modules employed for applications like cybersecurity/control monitoring and load forecasting in the SG environment, are often placed at a central server, they require the exchange of data between local Distributed Energy Resources (DERs) and the central node via communication networks. However, this presents privacy risks (sun2021solar) as the data being transferred can be utilized to reveal energy consumption patterns of private individuals (liu2018practical; zhao2020smart; wang2020spds). This is a major risk as the consumption patterns can be cross-analyzed to determine when the individuals are at home/outside. Further, their home appliances can also be revealed motivating malicious entities to schedule burglary attempts (liu2018practical).
@@
Another risk that the communication-based training data exchange between local DERs and the centralized server housing AI modules is the non-willingness within third-party DER owners to share proprietary data. This non-willingness typically stems from the concern that the proprietary data may be accessed by malicious entities who can then know the system details of the DER. This can be used to manipulate the system via stealth cyberattack vectors (leng2021stability). Non-willingness can also arise from a lack of trust between the DER owners and the utility company that manages the SG and the control center.
@@
Both the above-mentioned factors can lead to data constraints that may severely hinder the AI training process leading to flawed and untrustworthy decision-making, raising concerns on the reliability of AI. To solve the problem, An AI training strategy called Federated Learning can be used. In this strategy proprietary training data does not need to be transferred over a communication channel leading to alleviation of privacy concerns (mcmahan2017communication). More details about this training strategy are available in the following subsection.
@@ 2.2. Federated Learning
Federated Learning (FL) aims to collaboratively train models across multiple decentralized data owners holding potentially sensitive local data in a privacy-preserving manner. The essence of FL lies in its ability to learn a shared model by aggregating locally computed updates, rather than directly accessing or sharing the raw data. This machine learning (ML) paradigm not only enhances privacy and security but also enables the utilization of distributedly owned data for model training, making it particularly suitable for mission-critical applications, such as healthcare and finance, where data privacy is of paramount concern.
@@
To achieve meaningful real-world impact, FL must address several intricate challenges, including non-independently and identically distributed (non-IID) data across clients, system heterogeneity, and scalability. Non-IID data can introduce biases into models by favoring the data distribution of specific participants. System heterogeneity, characterized by variations in computation and communication capabilities among client participants, can result in uneven contributions to the model training process. Additionally, scalability concerns arise as the number of participants increases, necessitating the development of efficient algorithms and infrastructure to manage the aggregation of updates and distribution of the global model. The development of FL continues to evolve, spurred by advancements in both theoretical research and practical applications across various industries. FL is anticipated to play a pivotal role in building AI solutions that prioritize user privacy and data sovereignty.
The development of FL continues to evolve, spurred by advancements in both theoretical research and practical applications across various industries. FL is anticipated to play a pivotal role in building AI solutions that prioritize user privacy and data sovereignty.
@@
Consider a typical FL setup involving $m$ IoT devices (e.g., smart grid sensors) and a central smart grid server, where the devices collaborate to train a global model $\mathbf{w}\in\mathbb{R}^{p}$ with dimension $p$, coordinated by the central server. Assume each IoT device $j\in[m]$ holds a local private dataset $\mathcal{D}_{j}$. The goal of FL is to solve the following optimization problem:
(1) $\min_{\mathbf{w}\in\mathbb{R}^{p}}F(\mathbf{w}):=\frac{1}{m}\sum_{j\in[m]}F_{j}(\mathbf{w}),$
@@
where $F_{j}(\mathbf{w}):=\mathbb{E}_{z\in\mathcal{D}_{j}}[\ell(\mathbf{w};z)]$ represents the local loss for device $j$, and $\ell(\mathbf{w};z)$ is the loss function depending on the model parameter $\mathbf{w}$ and a data point $z$ sampled from $\mathcal{D}_{j}$. Note that the datasets $\mathcal{D}_{j}$ and $\mathcal{D}_{k}$ may have different distributions for $j\neq k$.
@@
To solve the above optimization problem, the system runs for $T$ rounds of FL training. Initially, the server stores the global model $\mathbf{w}^{0}$. In round $t$, the server randomly selects a subset $\mathcal{Q}^{t}$ of $q$ devices and sends them the latest global model $\mathbf{w}^{t}$. These selected devices update the global model $\mathbf{w}^{t}$ using their local data. For instance, in the widely-used Federated Averaging (FedAvg) (fl_original) algorithm, each selected device performs $\kappa$ iterations of stochastic gradient descent (SGD) to update the local model as follows:
(2) $\mathbf{w}_{j}^{t,r+1}=\mathbf{w}_{j}^{t,r}-\alpha\cdot\mathbf{g}_{j}^{t,r},\quad\forall r=0,\ldots,\kappa-1,$
@@
where $\mathbf{w}_{j}^{t,r}$ represents the local model of device $j$ during its training, initialized as $\mathbf{w}_{j}^{t,0}=\mathbf{w}^{t}$. Here, $\alpha$ is the local learning rate, and $\mathbf{g}_{j}^{t,r}:=\frac{1}{B}\sum_{z\in\xi_{j}^{t,r}}\nabla\ell(\mathbf{w}_{j}^{t,r},z)$ is the stochastic gradient computed over a mini-batch $\xi_{j}^{t,r}$ of size $B$ sampled from $\mathcal{D}_{j}$. After local training, each device computes the model update $\Delta_{j}^{t}:=\mathbf{w}^{t}-\mathbf{w}_{j}^{t,\kappa}$ and sends it to the server. In round $t$, the server aggregates the updates and applies them to the global model as follows:
(3) $\mathbf{w}^{t+1}=\mathbf{w}^{t}-\frac{1}{q}\sum_{j\in\mathcal{Q}^{t}}\Delta_{j}^{t},$
where $q=|\mathcal{Q}^{t}|$ is the number of selected IoT devices in round $t$.

### v3

@@ Abstract.
The Smart Grid (SG) is a critical energy infrastructure that collects real-time electricity usage data to forecast future energy demands using information and communication technologies (ICT). Due to growing concerns about data security and privacy in SGs, federated learning (FL) has emerged as a promising training framework. FL offers a balance between privacy, efficiency, and accuracy in SGs by enabling collaborative model training without sharing private data from IoT devices. In this survey, we thoroughly review recent advancements in designing FL-based SG systems across three stages: generation, transmission and distribution, and consumption. Additionally, we explore potential vulnerabilities that may arise when implementing FL in these stages. Furthermore, we discuss the gap between state-of-the-art (SOTA) FL research and its practical applications in SGs, and we propose future research directions. Unlike traditional surveys addressing security issues in centralized machine learning methods for SG systems, this survey is the first to specifically examine the applications and security concerns unique to FL-based SG systems. We also introduce FedGridShield, an open-source framework featuring implementations of SOTA attack and defense methods. Our aim is to inspire further research into applications and improvements in the robustness of FL-based SG systems.
@@
copyright: acmlicensedisbn: $c$s: Security and privacy Distributed systems security
@@ 1. Introduction
The Smart Grid (SG), which integrates the physical infrastructure of the electrical grid with modern information and communication technologies (ICTs) (Vasisht et al., 2022), is a prime example of a cyber-physical system (CPS). Unlike conventional power grids, SGs can detect, respond to, and proactively manage changes in electricity demand and supply, while also addressing various operational challenges (Obeagu and Owunna, 2022).
The SG architecture incorporates several key components that work together to enhance the overall functionality of the electrical system (Mittal et al., 2024), including advanced metering infrastructure (AMI) (Molokomme et al., 2019), communication networks (Hu et al., 2023b), and sophisticated information technology solutions (Reza et al., 2014), among others. By leveraging real-time data and advanced analytics, SGs aim to improve the efficiency of power generation, transmission and distribution, and consumption (Chopda et al., 2021). Although SGs enable a more flexible and resilient power system that can better accommodate the growing demand for electricity, the increasing reliance on ICTs has made SG systems more vulnerable to security and privacy attacks. Federated Learning (FL) has emerged as a promising approach to addressing privacy and security concerns in SG systems. This innovative distributed machine learning paradigm enables collaborative model training across multiple participants without the need to share raw data, making it particularly suitable for SG applications, where generators are distributed and data privacy for each generator is crucial. We categorize the application of FL in SGs into three aspects: generation, transmission and distribution, and consumption. From the perspective of generation (Tuballa and Abundo, 2016) in SGs, FL is primarily being used to enhance renewable energy integration and forecasting (Camacho et al., 2011). FL enables distributed SG generators to collaboratively train models for predicting resources such as solar and wind power output without sharing sensitive operational data. In the transmission (Bayoumi, 2015) and distribution aspects (Hamidi et al., 2010) of SGs, FL is proving valuable for line fault detection and classification, optimal power flow (Usman and Capitanescu, 2022), voltage control (Xie et al., 2021), and more. By utilizing an FL-based approach, grid operators can quickly identify and locate faults in transmission lines, while collaboratively training models for power distribution without compromising the privacy of individual generators. In the consumption process of SGs (Kaur et al., 2020), FL is extensively used for load forecasting (Husnoo et al., 2023) and energy theft detection (Li et al., 2024a). Smart meters can collaboratively train models to predict household energy consumption patterns without sharing individual usage data.
@@
Ref. Year Stage Applications DL-based Privacy Security FL-based
Grataloup et al. (Grataloup et al., 2024) 2024 $\surd$ $\surd$ $\surd$ $\surd$
Zheng et al. (Zheng et al., 2024) 2024 $\surd$ $\surd$ $\surd$ $\surd$
Zhang et al. (Zhang et al., 2024a) 2024 $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
Pandya et al. (Pandya et al., 2023) 2023 $\surd$ $\surd$ $\surd$ $\surd$
Cheng et al. (Cheng et al., 2022a) 2022 $\surd$ $\surd$ $\surd$ $\surd$
Mirzaee et al. (Mirzaee et al., 2022) 2022 $\surd$ $\surd$ $\surd$
Ding et al. (Ding et al., 2022) 2022 $\surd$ $\surd$ $\surd$ $\surd$
Cui et al. (Cui et al., 2020) 2020 $\surd$ $\surd$ $\surd$ $\surd$
@@ 1.1. Related Surveys
We summarize the state-of-the-art (SOTA) survey papers (Grataloup et al., 2024; Zheng et al., 2024; Zhang et al., 2024a; Pandya et al., 2023; Cheng et al., 2022a; Mirzaee et al., 2022; Ding et al., 2022; Cui et al., 2020; Gao et al., 2025) focus on SGs, with particular attention to seven aspects: whether the paper discusses the stages of SGs, compares SG applications, includes Deep Learning (DL)-based methods, discusses privacy and security issues, and whether it focuses on FL-based SG systems (see Table 1).
While existing surveys have explored various facets of FL or SG security, a dedicated and comprehensive analysis of FL-based SG systems—encompassing diverse applications alongside their specific security and privacy vulnerabilities—remains less developed. For instance, works by Grataloup et al. (Grataloup et al., 2024) and Pandya et al. (Pandya et al., 2023) touch upon FL applications in renewable energy and broader smart city contexts respectively, yet provide limited detail on SG-specific FL security issues. Other surveys, such as Zheng et al. (Zheng et al., 2024), detail FL components without addressing model-related attacks within FL-based SG systems, and Cheng et al. (Cheng et al., 2022a) offer a concise review of FL for SGs but lack depth on robustness. Conversely, while Zhang et al. (Zhang et al., 2024a), Mirzaee et al. (Mirzaee et al., 2022), Ding et al. (Ding et al., 2022), and Cui et al. (Cui et al., 2020) provide valuable insights into general SG security, common privacy violations, and specific threats like False Data Injection Attacks, their primary focus often remains on traditional centralized machine learning paradigms rather than the unique challenges introduced by FL. Our paper aims to bridge this crucial gap by providing a focused, comprehensive discussion on FL-based SGs. We systematically examine their applications across different SG stages and conduct a detailed analysis of their potential security and privacy vulnerabilities, thereby complementing existing research and intending to inspire further specialized studies at this critical intersection.
@@ 1.2. Key Contributions
    Our survey extensively discusses SOTA attack and defense methods in FL and identifies a significant gap between current research and its practical application in FL-based SG systems. To help bridge this gap and empower researchers to explore the SOTA strategies discussed, we introduce FedGridShield. This open-source framework provides implementations of these critical attack and defense methods, serving as a practical tool for pursuing the future research directions we further suggest in this domain.
@@ 1.3. Structure of the Paper
The structure of our paper is illustrated in Figure 1. In the following sections, we begin by exploring the foundational terminologies of SGs, the FL framework, and its threat models in Section 2. Next, Section 3 provides a comprehensive review of FL-based SG applications, detailing their implementation across various stages. In Section 4, we thoroughly examine vulnerabilities within SG applications. This analysis covers traditional adversarial methods as well as SOTA FL-based attack and defense methods relevant to these systems. Within this section, we also identify existing research gaps and propose future directions in the field. Subsequently, in Section 5, we introduce our FedGridShield framework and detail its implementation. Finally, Section 6 concludes the paper.
@@ 2.1. Smart Grid
The electric power grid has evolved by leaps and bounds since its inception (Masera et al., 2018). A major development in this domain was the integration of ICT. The integration of ICT with the traditional power grid facilitated distributed power generation via better resource management and automated control (Simões et al., 2012). ICT integration also marked the evolution of the conventional grid into the SG and solved several limitations, including unidirectional power flow and centralized bulk generation of electricity. Additionally, it reduced the power and energy domain’s dependence on fossil fuels like coal and natural gas, leading the way for sustainability and decarbonization via the development of microgrids (Jahromi et al., 2023; Yadav et al., 2020). The SG power flow framework consists of the following stages:
@@
    Generation: This stage refers to the production of AC/DC power. In the conventional power grid, power generation was centralized in nature. However, with ICT integration for real-time monitoring, power generation can also be achieved in a regulated manner at residential buildings (for example, using solar photovoltaic panels) and used for local consumption/fed back to the grid (Kahrobaee et al., 2012).
@@
    Transmission/Distribution: This stage refers to the transfer of the power produced at the generation stage to the appropriate end nodes for utilization. This involves the use of transformers, transmission lines, substation equipment, and other line devices like fault detector relays, breakers to ensure a safe and reliable transfer at appropriate bus voltage levels. The fault detector relays and breakers are often digital in nature and controlled by the central monitoring station via ICT devices (Gaspar et al., 2023).
@@
    Consumption: This is the final stage of the SG power flow process and involves the utilization of the power as received from the transmission and distribution networks for powering electrical devices. The consumption level can involve entities like government agencies, private industries, private individuals, etc., each of which can have a different consumption footprint. In the modern SG, integration of ICT devices means that some consumer-level devices like electric vehicles can also be used to transfer surplus/unused power back to the grid (coordinated via V2G communications (Madawala and Thrimawithana, 2011)).
@@
While ICT integration in the power grid enabled easier and more robust resource management from a single centralized Supervisory Control and Data Acquisition (SCADA)-based control system (Gaushell and Block, 1993), the increasing number of IoT devices and local generation facilities like solar panels means that it can be extremely computationally resource-intensive to monitor them. Inadequate monitoring facilities can mean that intrusive events like cyberattacks and line faults may not be detected and mitigated on time.
@@
Slowed mitigation of cyberattacks, faults, and other control emergencies can have a devastating impact on the stability of the SG (Jahromi et al., 2023). Moreover, the SG being a cyber-physical system implies that cyber-layer vulnerabilities can be exploited to cause severe physical ramifications. For example, cyberattacks on the SG can lead to uncontrolled voltage increase at one or more buses leading to equipment damage and consumer-level device burnout (Rath et al., 2022). To alleviate these issues, several researchers have proposed, implemented, and evaluated the use of Artificial Intelligence (AI)-based tools and techniques for applications like real-time control (Bose, 2017), cybersecurity (Ruan et al., 2023), load forecasting (Zhou and Jin, 2017), and fault detection (Sun et al., 2023). These AI tools are typically deployed at the centralized controller level for monitoring, removing the need for human system operators.
@@
As the AI modules employed for applications like cybersecurity/control monitoring and load forecasting in the SG environment, are often placed at a central server, they require the exchange of data between local Distributed Energy Resources (DERs) and the central node via communication networks. However, this presents privacy risks (Sun et al., 2021) as the data being transferred can be utilized to reveal energy consumption patterns of private individuals (Liu et al., 2018). This is a major risk as the consumption patterns can be cross-analyzed to determine when the individuals are at home/outside. Further, their home appliances can also be revealed motivating malicious entities to schedule burglary attempts (Liu et al., 2018).
@@
Another risk of the communication-based training data exchange between local DERs and the centralized server housing AI modules is the non-willingness within third-party DER owners to share proprietary data. This non-willingness typically stems from the concern that the proprietary data may be accessed by malicious entities who can then know the system details of the DER. This can be used to manipulate the system via stealth cyberattack vectors (Leng et al., 2021). Non-willingness can also arise from a lack of trust between the DER owners and the utility company that manages the SG and the control center.
@@
Both the above-mentioned factors can lead to data constraints that may severely hinder the AI training process, leading to flawed and untrustworthy decision-making, raising concerns on the reliability of AI. To solve the problem, an AI training strategy called FL can be used. In this strategy, proprietary training data does not need to be transferred over a communication channel, leading to alleviation of privacy concerns (McMahan et al., 2017a). More details about this training strategy are available in the following subsection.
@@ 2.2. Federated Learning
FL (Zhang et al., 2024b, 2025) aims to collaboratively train models across multiple decentralized data owners holding potentially sensitive local data in a privacy-preserving manner. The essence of FL lies in its ability to learn a shared model by aggregating locally computed updates, rather than directly accessing or sharing the raw data. This machine learning (ML) paradigm not only enhances privacy and security but also enables the utilization of distributively owned data for model training, making it particularly suitable for mission-critical applications, such as healthcare and finance, where data privacy is of paramount concern.
@@
To achieve significant real-world impact, FL must overcome several complex challenges. These include the prevalence of non-independently and identically distributed (non-IID) data across clients, systemic heterogeneity, and scalability issues. Non-IID data can introduce biases, favoring the data distributions of particular participants. System heterogeneity, arising from variations in computational power and communication capabilities among clients, can lead to uneven contributions to the model training process. Furthermore, as the number of participants grows, scalability becomes a concern, necessitating efficient algorithms and robust infrastructure for managing update aggregation and global model distribution. The field of FL is continuously evolving, driven by advancements in both theoretical research and practical applications across diverse industries, and is anticipated to play a pivotal role in developing AI solutions that prioritize user privacy and data sovereignty.
@@
Consider a typical FL setup involving $m$ clients (e.g., IoT devices such as smart grid sensors) and a central server (e.g., a smart grid server). These clients collaborate to train a global model $\mathbf{w}\in\mathbb{R}^{p}$, with dimension $p$, coordinated by the central server. Assume each client $j\in[m]$ possesses a local private dataset $\mathcal{D}_{j}$. The primary objective of FL is to solve the following optimization problem:
(1) $\min_{\mathbf{w}\in\mathbb{R}^{p}}F(\mathbf{w}):=\frac{1}{m}\sum_{j\in[m]}F_{j}(\mathbf{w})$
@@
where $F_{j}(\mathbf{w}):=\mathbb{E}_{z\in\mathcal{D}_{j}}[\ell(\mathbf{w};z)]$ represents the local loss function for client $j$, and $\ell(\mathbf{w};z)$ is the loss function dependent on the model parameter $\mathbf{w}$ and a data point $z$ sampled from $\mathcal{D}_{j}$. It is important to note that the datasets $\mathcal{D}_{j}$ and $\mathcal{D}_{k}$ may exhibit different distributions for $j\neq k$.
@@
To solve this optimization problem, the system operates for $T$ rounds of FL training. Initially, the server stores the global model $\mathbf{w}^{0}$. In each round $t\in[T]$, the server randomly selects a subset $\mathcal{Q}^{t}$ of $q$ clients and disseminates the latest global model $\mathbf{w}^{t}$ to them. These selected clients then update the global model $\mathbf{w}^{t}$ using their respective local data. For instance, in the widely-used FedAvg algorithm, each selected client $j\in\mathcal{Q}^{t}$ performs $\kappa$ iterations of stochastic gradient descent (SGD) to update its local model as follows:
(2) $\mathbf{w}_{j}^{t,r+1}=\mathbf{w}_{j}^{t,r}-\alpha\cdot\mathbf{g}_{j}^{t,r},\quad\forall r=0,\ldots,\kappa-1$
@@
where $\mathbf{w}_{j}^{t,r}$ denotes the local model of client $j$ at local iteration $r$ of round $t$, initialized as $\mathbf{w}_{j}^{t,0}=\mathbf{w}^{t}$. Here, $\alpha$ is the local learning rate, and $\mathbf{g}_{j}^{t,r}:=\frac{1}{B}\sum_{z\in\xi_{j}^{t,r}}\nabla\ell(\mathbf{w}_{j}^{t,r},z)$ is the stochastic gradient computed over a mini-batch $\xi_{j}^{t,r}$ of size $B$ sampled from $\mathcal{D}_{j}$. After local training, each client computes the model update $\Delta_{j}^{t}:=\mathbf{w}^{t}-\mathbf{w}_{j}^{t,\kappa}$ and sends it to the server. In round $t$, the server aggregates these updates and applies them to the global model. However, the presence of malicious actors necessitates considering potential attacks.
2.3. Threat Models
We consider different types of adversaries who can compromise the integrity and privacy of the FL system. Let $\mathcal{A}$ denote a generic attacker.
Privacy Inference Attackers. Let $\mathcal{A}_{priv}$ denote a privacy inference attacker. The objective of $\mathcal{A}_{priv}$ is to deduce sensitive information about clients’ local data $\mathcal{D}_{j}$ by observing the shared model updates $\Delta_{j}^{t}$ or the aggregated global model $\mathbf{w}^{t}$. Their capabilities ($\mathcal{C}_{priv}$) includes,
  • Possess diverse prior information (e.g., auxiliary datasets, knowledge of model architecture) and computational resources.
  • May collude with other malicious entities (clients or an adversarial server).
  • Typically do not inject fabricated data with the intent of degrading the model, but passively observe communications to infer private data.
  • Can be an “honest-but-curious” server or participating clients.
Byzantine Attackers. Let $\mathcal{A}_{byz}$ denote a Byzantine attacker. These attackers aim to disrupt the learning process or degrade the performance of the global model. The objective of a general Byzantine attacker, $\mathcal{A}_{byz}$, within the FL system is to undermine the overall performance of the trained global model $\mathbf{w}$ as much as possible. This is also known as an untargeted model poisoning attack. Let $k_{byz}$ be the number of Byzantine attackers among the $m$ clients, where typically $k_{byz}<m/2$ for many defense strategies to be effective. Their capabilities ($\mathcal{C}_{byz}$) often includes:
  • Maliciously alter their local SGD processes (Equation (2)), for example, by flipping gradients, adding large noise to updates, or submitting random model weights.
  • May generate and submit arbitrary model updates $\tilde{\Delta}_{byz}^{t}$ that do not correspond to legitimate training on their local data.
  • May collude with other Byzantine attackers to coordinate their malicious updates.
  • Due to secure aggregation, they typically remain unaware of the model updates $\Delta_{j}^{t}$ of benign clients in the same round before aggregation.
Backdoor Attackers. Let $\mathcal{A}_{bd}$ denote a backdoor attacker. This is a specific type of model poisoning attacker with a more targeted malicious goal. The objective of $\mathcal{A}_{bd}$ is to implant a hidden backdoor into the global model $\mathbf{w}$. The backdoored model $\mathbf{w}_{bd}$ should behave normally on standard inputs from the primary task distribution but produce specific, attacker-chosen incorrect outputs when inputs contain a predefined trigger pattern (e.g., a small pixel patch in an image, a specific phrase in text). Their capabilities ($\mathcal{C}_{bd}$) includes:
  • Control a fraction of the participating clients (let this be $k_{bd}$ clients).
  • Modify the training data $\mathcal{D}_{j}$ on their controlled clients by injecting samples that pair the backdoor trigger with the target malicious label.
  • Manipulate the local model training process (Equation (2)) on their controlled clients to generate model updates $\Delta_{j}^{t,bd}$ that steer the global model towards learning the backdoor task.
  • Aim for the backdoor to be persistent and effective in the aggregated global model $\mathbf{w}$ without significantly degrading its performance on the main task for benign inputs (to evade detection).
  • May possess knowledge of the model architecture and the aggregation scheme.